Find the number of all the ordered arrangements of 6 objects a, b, c, d, e, and f, choosing 4 at a time without repetition.

Answer by jim_thompson5910(35256)

Since order matters, this means that there are 6 P 4 = 6!/(6-4)! = 6!/2! = 720/2 = 360 different arrangements.


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim